# Logos as a Well-Tempered Pre-train for Sign Language Recognition

### v1

@@ Abstract
This paper examines two aspects of the isolated sign language recognition (ISLR) task. First, despite the availability of a number of datasets, the amount of data for most individual sign languages is limited. It poses the challenge of cross-language ISLR model training, including transfer learning. Second, similar signs can have different semantic meanings. It leads to ambiguity in dataset labeling and raises the question of the best policy for annotating such signs. To address these issues, this study presents Logos, a novel Russian Sign Language (RSL) dataset, the most extensive ISLR dataset by the number of signers and one of the largest available datasets while also the largest RSL dataset in size and vocabulary. It is shown that a model, pre-trained on the Logos dataset can be used as a universal encoder for other language SLR tasks, including few-shot learning. We explore cross-language transfer learning approaches and find that joint training using multiple classification heads benefits accuracy for the target low-resource datasets the most. The key feature of the Logos dataset is explicitly annotated visually similar sign groups. We show that explicitly labeling visually similar signs improves trained model quality as a visual encoder for downstream tasks. Based on the proposed contributions, we outperform current state-of-the-art results for the WLASL dataset and get competitive results for the AUTSL dataset, with a single stream model processing solely RGB video. The source code, dataset, and pre-trained models are publicly available.
(a) Летать (to fly) Refer to caption
(b) Крыло (wing) Refer to caption
(c) Овца (sheep) Refer to caption
@@ 1 Introduction
Sign languages (SL) are visual-spatial signals for communication among deaf communities. Primarily, information in these languages is expressed by hand shape and their motion (manual components), but also with a great aid of motion of mouth, head, eyes, and the body (non-manual components).
@@
As a rule, a national sign language is associated with a national spoken language. Each sign corresponds to a spoken word named “gloss”, which describes the sign’s meaning. However, sign languages are independent languages with their own vocabulary and rules Sandler (2006).
The presented work deals with the isolated sign language recognition (ISLR) problem, i.e., the classification of videos that contain only one sign each. The ISLR task has not only independent significance but is also important for building a more practical continuous sign language translation (CSLT) solution Chen et al. (2022); Wei and Chen (2023); Zuo et al. (2024).
The serious obstacle to building SLR solutions is a shortage of training data Gokul et al. (2022); Papadimitriou and Potamianos (2023). While a number of annotated SL datasets exist, they represent different sign languages (Table 1), and dataset corpora for many individual languages are insufficient. It highlights the task of cross-lingual use of data. Although some papers utilize cross-lingual training, they are limited to joint use of two or more rather small datasets. To our knowledge, no research investigates the ability of a model trained on an extensive SL dataset to serve as an encoder for SL tasks for other sign languages and compares different approaches to it. While the applicability of a larger dataset of the same language is straightforward, cross-language transfer learning needs an investigation. This paper presents an extensive Russian Sign Language (RSL) dataset, Logos, one of the largest existing sign language datasets in terms of volume and vocabulary size and the largest in terms of the number of signers. We show that a model pre-trained on the Logos dataset can be successfully transferred to another language SLR tasks, including few-shot learning. The dataset size is critical, and the effect degrades if a smaller dataset is used for pre-train. Next, we compare transfer learning methods and find that simultaneous training with the large dataset using multiple classification heads for different languages benefits the target language SLR models the most, compared to other transfer learning methods.
Another problem with SLR is that signs with similar handshapes and motions can have various semantic meanings. Such signs can be either strictly indistinguishable or only distinguishable by their constituent non-manual features Zuo et al. (2023); Hu et al. (2021b), see Figure 1. The difference between the individual signers’ manner blurs the boundary between non-manual features and makes such signs practically indistinguishable out of context. This paper calls such hardly distinguishable signs “visually similar signs” (VSSigns).
@@
  • We show that a model, pre-trained on the proposed Logos dataset can be transferred to another language SLR tasks, including few-shot learning. We compare transfer learning methods and demonstrate that the method of cross-lingual multi-dataset co-training with multiple language-specific classification heads improves SL models for low-resource datasets the most, compared to the conventional “pre-train and finetune” method.
@@
The research was conducted in cooperation with the “All-Russian Society of the Deaf” (VOG). VOG experts and professional sign language interpreters participated at every stage of the Logos dataset creation. We also engaged deaf consultants in developing training strategies to apply considerations to specific solutions. Additionally, some of our researchers completed formal courses on RSL to enhance their knowledge in this domain.
@@ 2 Related Works
DEVISIGN-L Wang et al. (2016) lab Chinese (CSL) 24,000 8 2,000
SLR500 Huang et al. (2018) lab Chinese (CSL) 125,000 50 500
MS-ASL Joze and Koller (2018) web American (ASL) 25,513 222 1,000 grouped
SMILE Ebling et al. (2018) lab Swiss German (DSGS) 9,000 30 100
BosphorusSign22k Özdemir et al. (2020) lab Turkish (TSL) 22,542 6 744 grouped
AUTSL Sincan and Keles (2020) lab Turkish (TSL) 38,336 43 226
WLASL Li et al. (2020) web American (ASL) 21,083 119 2,000
BSLDict Momeni et al. (2020) lab British (BSL) 14,210 28 9,283 addressed
BSL-1K Albanie et al. (2020) TV British (BSL) 273,000 40 1,064
INCLUDE Sridhar et al. (2020) lab Indian (ISL) 4,292 7 263
NMFs-CSL Hu et al. (2021b) lab Chinese (CSL) 32,010 10 1,067 addressed
BOBSL Albanie et al. (2021) TV British (BSL) 452,000 39 2,281
GSL isol. Adaloglou et al. (2021) lab Greek (GSL) 40,785 7 310 grouped
LSFB-ISOL Fink et al. (2021) lab Fra/Bel 47,600 100 395
CISLR Joshi et al. (2022) web Indian (ISL) 7,000 71 4,765
LSA64 Ronchetti et al. (2023) lab Argentinian 3,200 10 64
ASL Citizen Desai et al. (2024) crowd American (ASL) 83,399 52 2,731
Slovo Kapitanov et al. (2023) crowd Russian (RSL) 20,000 194 1,000
FDMSE-ISL Patra et al. (2024) lab Indian (ISL) 40,033 20 2,000
MM-WLAuslan Shen et al. (2024b) lab Australian(Auslan) 282,000 76 3,215
Logos (Ours) crowd Russian (RSL) 200,000 381 2,863/2,004∗∗ both
Table 1: Summary of existing ISLR datasets. Method – the collection method: laboratory, web scrapping, TV programs, crowdsourcing. VSSigns column shows if visually similar signs (VSSigns) were considered by the dataset authors: grouped – VSSigns groups were assigned a unique label; addressed – the authors adopt VSSigns presence in the dataset and propose some methods to tackle them at training time; the dash – VSSigns presence is not discussed. — these datasets mostly have automatic annotations of isolated glosses. ∗∗ — numbers of ungrouped gloss labels and grouped VSSign labels are provided.
@@ 2.1 Isolated Sign Language Recognition
In recent years, a group of approaches for ISLR tasks rely on using RGB input data. Then, either 2D convolutional neural network (CNN) is applied to extract individual frames’ features, followed by LSTM for the temporal component processing Koller et al. (2019), or the spatial and temporal components are simultaneously processed using 3D CNN Papadimitriou and Potamianos (2023); Zuo et al. (2023); Albanie et al. (2020); Huang et al. (2018); Li et al. (2020); Joze and Koller (2018). After the proliferation of transformers, transformer-based image and video processing architectures were applied Kapitanov et al. (2023); Kvanchiani et al. (2024). In addition to the RGB input, a depth map can be used Jiang et al. (2021); Zuo et al. (2023).
Another group of approaches utilizes pose (skeleton) keypoints and face landmarks generated by available frameworks Hrúz et al. (2022); Jiang et al. (2021); Miah et al. (2023); Papadimitriou and Potamianos (2023); Ryumin et al. (2023). The skeleton keypoints can be represented as a sequence of heatmaps and processed similarly to video data Zuo et al. (2023). A series of methods build a graph based on physical skeleton connections and explore Graph Convolutional Networks (GCNs) Hu et al. (2021a, 2023); Patra et al. (2024); Zhao et al. (2023); Jiang et al. (2021).
Most current SOTA SLR models are multi-stream and multi-modal and combine more than one of the methods listed above Hrúz et al. (2022); Zuo et al. (2023); Jiang et al. (2021); Miah et al. (2023); Papadimitriou and Potamianos (2023); Ryumin et al. (2023).
@@ 2.2 ISLR Datasets
The ISLR datasets differ in several aspects: language, collection method, size, vocabulary size, number of signers (see Table 1). The most common method of dataset collection is recording invited seiners in laboratory conditions. However, this approach generally results in insufficient scene and signer variety, requiring the authors to record each video individually. Web scrapping of SL videos is rather effective and results in more diverse datasets. However, its serious problem is the absence of consent from the video owner and person represented in the video on the usage of the video as a part of the dataset. Albanie et al. Albanie et al. (2020, 2021) prepared the British SL datasets using BBC TV programs with SL translation. The datasets are large but have limited scene variety and number of signers, and they mostly only have automatic annotation. Collecting video from SL experts using a web crowdsourcing platform has no problem with signers’ consent and provides much more diverse footage. We have used this approach for our work.
@@ 2.3 The VSSigns Problem
There is no standard approach to annotating visually similar signs (VSSigns). As a result, similar signs for different glosses can be annotated with either different or similar labels. In this paper, we call it ungrouped gloss and grouped VSSign annotations. The datasets collected for the most common words of spoken language Sincan and Keles (2020); Kapitanov et al. (2023), typical continuous phrases Albanie et al. (2020, 2021); Adaloglou et al. (2021), or based on an SL dictionary Patra et al. (2024) primarily have different (ungrouped) labels for similar signs. For instance, according to Zuo et al. (2023), among 2,000 classes of widely used WLASL dataset Li et al. (2020), 334 classes form groups of VSSigns. Additional efforts are needed to merge similar VSSigns and assign unique grouped VSSign labels to them.
Among the reviewed datasets, three papers state that VSSigns were grouped. Two papers confirm the presence of ungrouped VSSigns in the presented datasets and propose some techniques to distinguish them (Table 1). To improve VSSigns classification, Hu et al. Hu et al. (2021b) deform a feature map, stretching more informative areas to emphasize non-manual features. Zuo et al. Zuo et al. (2023) propose label smoothing depending on their semantic difference and a common latent space for gloss embeddings and vision features to maximize the separability of confusing signs. Other works do not mention any steps to handle VSSigns in the proposed datasets. To our knowledge, no research has examined the impact of VSSigns on the quality of the resulting encoder for downstream tasks. Such a study is one of the topics of this work, using transfer learning to another language as a downstream task example.
@@ 2.4 Multi-Dataset Training
Although researchers complain about insufficient SLR training data Gokul et al. (2022); Papadimitriou and Potamianos (2023), the topic of cross-language dataset sharing is poorly exploited. Gokul et al. Gokul et al. (2022) implemented a multilingual SLR model for 11 sign languages by simply translating the labels of all the languages into English. The authors themselves admit that their model of combining different languages is primitive and does not make progress for some datasets. Tornay et al. Tornay et al. (2020) train a unified hand movement model using 3 different sign language resources. Then, they optimize the classifier using the target sign language data. However, their cross-lingual model falls short of the monolingual reference. Yin et al. Yin et al. (2022) propose the MLSLT translation network as a single model for multilingual translation. Their work is limited to their rather small datasets and doesn’t address leveraging large SL datasets to improve the model quality. Hu et al. Hu et al. (2022) introduced an additional shared module that learns knowledge from two languages. It improved accuracy for Chinese CSL-Daily Zhou et al. (2021) and German Phoenix-14 Koller et al. (2015) datasets. Wei et al. Wei and Chen (2023) also benefit from the joint using the same datasets by creating a gloss translation map based on the visual similarity of signs, rather than their meanings. Authors train the model for the German language using both datasets and replace gloss labels in Chinese videos with German labels using this map, treating Chinese signs as German. As shown below, this mapping method does not give optimal results (see Section 5.3).
However, no research has been found that investigates the ability of a model trained on an extensive SL dataset to serve as an encoder for SL tasks for other sign languages and compares different approaches to it. Such a study is another subject of our work.
Figure 2: Dataset characteristics and distribution analysis. a) Sign length distribution. b) Distance distribution. The distance (in meters) is approximately estimated based on the length between the left and right shoulders of the signer obtained using MediaPipe Lugaresi et al. (2019). c) Signers’ devices. d) Devices resolution. e) Number of videos per signer. f) Brightness distribution. The sample brightness is the mean pixel brightness of grayscaled video frames. g) Signers’ gender; h) Signers’ age. The age is determined by the MiVOLO model Kuprashevich and Tolstykh (2023).

### v2

@@ Abstract
This paper examines two aspects of the isolated sign language recognition (ISLR) task. First, although a certain number of datasets is available, the data for individual sign languages is limited. It poses the challenge of cross-language ISLR model training, including transfer learning. Second, similar signs can have different semantic meanings. It leads to ambiguity in dataset labeling and raises the question of the best policy for annotating such signs. To address these issues, this study presents Logos, a novel Russian Sign Language (RSL) dataset, the most extensive available ISLR dataset by the number of signers, one of the most extensive datasets in size and vocabulary, and the largest RSL dataset. It is shown that a model, pre-trained on the Logos dataset can be used as a universal encoder for other language SLR tasks, including few-shot learning. We explore cross-language transfer learning approaches and find that joint training using multiple classification heads benefits accuracy for the target low-resource datasets the most. The key feature of the Logos dataset is explicitly annotated visually similar sign groups. We show that explicitly labeling visually similar signs improves trained model quality as a visual encoder for downstream tasks. Based on the proposed contributions, we outperform current state-of-the-art results for the WLASL dataset and get competitive results for the AUTSL dataset, with a single stream model processing solely RGB video. The source code, dataset, and pre-trained models are publicly available.
Logos as a Well-Tempered Pre-train for Sign Language Recognition
Ilya Ovodov, Petr Surovtsev, Karina Kvanchiani, Alexander Kapitanov, Alexander Nagaev SberAI {iovodov, petr.surovcev, karinakvanciani, kapitanovalexander, sashanagaev1111}@gmail.com
(a) Летать (to fly)
(b) Крыло (wing)
(c) Овца (sheep)
@@ 1 Introduction
Sign languages (SL) are visual-spatial signals for communication among deaf communities. Although national sign languages are mostly associated with national spoken languages, they are distinct languages with their own grammar and vocabulary. Primarily, signs are expressed by hand shape and motion (manual components of sign), but also with a great aid of motion of mouth, head, eyes, and the body (non-manual components). The term ‘‘gloss’’ is used to refer to the word that signifies the sign. Generally, this word represents the sign’s meaning. Still, one sign can be translated by several words and vice versa, so glosses should be considered just as word labels for signs.
@@
The presented work deals with the isolated sign language recognition (ISLR) problem, i.e., the classification of videos that contain only one sign each. The ISLR task has not only independent significance but is also important for building a more practical continuous sign language translation (CSLT) solution chen2022simple; wei2023ImprovingContinuousSign; zuo2024towards.
A significant obstacle to building SLR solutions is a shortage of training data gokul2022AddressingResourceScarcity; papadimitriou2023SignLanguageRecognition. While a number of annotated SL datasets exist, they represent different sign languages (Table 1), and dataset corpora for many individual languages are insufficient. It highlights the task of using cross-lingual data. Some researchers utilize cross-lingual training either by combining two or more rather small datasets or using a multilingual dataset. The presented study examines the ability of a model trained on an extensive SL dataset of one language to serve as an encoder for SL tasks for other sign languages, and compares different approaches to it.
This paper presents an extensive Russian Sign Language (RSL) dataset, Logos, one of the largest existing sign language datasets in terms of volume and vocabulary size and the largest in terms of the number of signers. We show that a model pre-trained on the Logos dataset can be successfully transferred to another language SLR tasks, including few-shot learning. The dataset size is critical, and the effect degrades if a smaller dataset is used for pre-train. Next, we compare transfer learning methods and find that simultaneous training with the large dataset using multiple classification heads for different languages benefits the target language SLR models the most, compared to other transfer learning methods.
Another problem with SLR is that signs with similar handshapes and motions can have various semantic meanings. Such signs can be either strictly indistinguishable (the cases of polysemy or homonymy) or only distinguishable by their constituent non-manual features zuo2023NaturalLanguageAssistedSign; hu2021GlobalLocalEnhancementNetwork, see Figure 1. In the latter case, such signs can be viewed as minimal pairs for SL. The difference between the individual signers’ manners blurs the boundary between non-manual features and makes such signs practically indistinguishable out of context. For this reason, while some researchers consider such signs to be separate, others consider them to be the same ebling2015factors. This paper calls such hardly distinguishable signs ‘‘visually similar signs’’ (VSSigns). This concept includes both polysemy and minimal pair cases. While they differ from a linguistic perspective, it makes sense to combine them for machine learning purposes.
@@
  • We show that a model, pre-trained on the proposed Logos dataset can be transferred to another language SLR tasks, including few-shot learning. We compare transfer learning methods and demonstrate that the method of cross-lingual multi-dataset co-training with multiple language-specific classification heads improves SL models for low-resource datasets the most, compared to the conventional ‘‘pre-train and finetune’’ method.
@@
The research was conducted in cooperation with the ‘‘All-Russian Society of the Deaf’’ (VOG). VOG experts and professional sign language interpreters participated at every stage of the Logos dataset creation. We also engaged deaf consultants in developing training strategies to apply considerations to specific solutions. Additionally, some of our researchers completed formal courses on RSL to enhance their knowledge in this domain.
@@ 2 Related Works
DEVISIGN-L wang2016IsolatedSignLanguage lab Chinese (CSL) 24,000 8 2,000
SLR500 huang2019AttentionBased3DCNNsLargeVocabulary lab Chinese (CSL) 125,000 50 500
MS-ASL joze2019MSASL web American (ASL) 25,513 222 1,000 grouped
SMILE ebling2018SMILESwissGerman lab Swiss German 9,000 30 100
BosphorusSign22k ozdemir2020BosphorusSign22kSignLanguage lab Turkish (TİD) 22,542 6 744 grouped
AUTSL sincan2020AUTSLLargeScale lab Turkish (TİD) 38,336 43 226
WLASL li2020WLASL web American (ASL) 21,083 119 2,000
BSLDict momeni2020WatchReadLookup lab British (BSL) 14,210 28 9,283 addressed
K-RSL imashev2020k lab Kazakh-Russian 28,250 10 600 addressed
BSL-1K albanie2020BSL1K TV British (BSL) 273,0001 40 1,064
INCLUDE sridhar2020INCLUDE lab Indian (ISL) 4,292 7 263
NMFs-CSL hu2021GlobalLocalEnhancementNetwork lab Chinese (CSL) 32,010 10 1,067 addressed
BOBSL albanie2021BBCOxfordBritishSign TV British (BSL) 452,0001 39 2,281
GSL isol. adaloglou2021ComprehensiveStudyDeep lab Greek (GSL) 40,785 7 310 grouped
LSFB-ISOL fink2021LSFB lab Fra/Bel 47,600 100 395
CISLR joshi2022CISLRCorpusIndian web Indian (ISL) 7,000 71 4,765
LSA64 ronchetti2023LSA64ArgentinianSign lab Argentinian 3,200 10 64
ASL Citizen desai2024ASLCitizenCommunitysourced crowd American (ASL) 83,399 52 2,731
Slovo kapitanov2023SlovoRussianSign crowd Russian (RSL) 20,000 194 1,000
FDMSE-ISL patra2024HierarchicalWindowedGraph lab Indian (ISL) 40,033 20 2,000
MM-WLAuslan shen2024MMWLAuslanMultiViewMultiModal lab Australian(Auslan) 282,0002 76 3,215
Logos (Ours) crowd Russian (RSL) 200,000 381 2,8633 both
Table 1: Summary of existing ISLR datasets. Method – the collection method: laboratory, web scrapping, TV programs, crowdsourcing. VSSigns column shows if visually similar signs (VSSigns) were considered by the dataset authors: grouped – VSSigns groups have common labels; addressed – the authors adopt VSSigns presence in the dataset and propose some methods to tackle them at training time; "–" – VSSigns presence is not discussed.
1 These datasets mostly have automatic annotations of isolated glosses.
2 Actually, the dataset contain 70730 samples recorderd from four views each.
3 This number of glosses is grouped into 2,004 VSSign labels.
@@ 2.1 Isolated Sign Language Recognition
In recent years, a group of approaches for ISLR tasks rely on using RGB input data. Then, either 2D convolutional neural network (CNN) is applied to extract individual frames’ features, followed by LSTM for the temporal component processing koller2019WeaklySupervisedLearning, or the spatial and temporal components are simultaneously processed using 3D CNN papadimitriou2023SignLanguageRecognition; zuo2023NaturalLanguageAssistedSign; albanie2020BSL1K; huang2019AttentionBased3DCNNsLargeVocabulary; li2020WLASL; joze2019MSASL. After the proliferation of transformers, transformer-based image and video processing architectures were applied kapitanov2023SlovoRussianSign; kvanchiani2024TrainingStrategiesIsolated. In addition to the RGB input, a depth map can be used jiang2021SkeletonAwareMultimodal; zuo2023NaturalLanguageAssistedSign.
Another group of approaches utilizes pose (skeleton) keypoints and face landmarks generated by available frameworks hruz2022OneModelNot; jiang2021SkeletonAwareMultimodal; miah2023MultistreamGeneralGraphbased; papadimitriou2023SignLanguageRecognition; ryumin2023AudiovisualSpeechGesture. The skeleton keypoints can be represented as a sequence of heatmaps and processed similarly to video data zuo2023NaturalLanguageAssistedSign. A series of methods build a graph based on physical skeleton connections and explore Graph Convolutional Networks (GCNs) hu2021SignBERT; hu2023signbert; patra2024HierarchicalWindowedGraph; zhao2023BESTBERTPretraining; jiang2021SkeletonAwareMultimodal.
Most current SOTA SLR models are multi-stream and multi-modal and combine more than one of the methods listed above hruz2022OneModelNot; zuo2023NaturalLanguageAssistedSign; jiang2021SkeletonAwareMultimodal; miah2023MultistreamGeneralGraphbased; papadimitriou2023SignLanguageRecognition; ryumin2023AudiovisualSpeechGesture.
Nevertheless, we focus our research on purely frontal RGB video, as this setup is most relevant for practical applications. Many possible real-world scenarios, such as educational tools for RSL and ASL learning, video conferencing, and public kiosk systems in environements like metro stations or airports, typically involve users facing a single RGB camera.
@@ 2.2 ISLR Datasets
The ISLR datasets differ in several aspects: language, collection method, size, vocabulary size, number of signers (see Table 1). The most common method of dataset collection is recording invited signers in laboratory conditions. However, this approach generally results in insufficient scene and signer variety, requiring the authors to record each video individually. Web scrapping of SL videos is rather effective and results in more diverse datasets. However, its serious problem is the absence of consent from the video owner and person represented in the video on the usage of the video as a part of the dataset. Albanie et al. albanie2020BSL1K; albanie2021BBCOxfordBritishSign prepared the British SL datasets using BBC TV programs with SL translation. The datasets are large but have limited scene variety and number of signers, and they mostly only have automatic annotation. Collecting video from SL experts using a web crowdsourcing platform has no problem with signers’ consent and provides much more diverse footage. We have used this approach for our work.
@@ 2.3 The VSSigns Problem
Some SL signs with different semantic meanings either can be considered as strictly indistinguishable (the cases of polysemy or homonymy) or can have similar handshapes and motions, but can only be distinguished by their constituent non-manual features zuo2023NaturalLanguageAssistedSign; hu2021GlobalLocalEnhancementNetwork, see Figure 1. The latter can be viewed as minimal pairs for SL. The border between polysemy and minimal pair signs can be blurred due to the individual signers’ manners, making even signs with different non-manual features practically indistinguishable out of context. For this reason, while some researchers consider such signs to be separate, others consider them to be the same ebling2015factors. On the contrary, even the same sign can have distinct manual features depending on the context, i.e., a question vs. a statement mukushev2020evaluation. It results in ambiguity in the annotation of such signs. This paper calls such hardly distinguishable signs ‘‘visually similar signs’’ (VSSigns). Formally, we define them as signs that have different meanings but have the same manual component. The VSSign concept includes both polysemy and minimal pair cases. While they differ from a linguistic perspective, it makes sense to combine them for machine learning purposes.
There is no standard approach to annotating VSSigns: they can be annotated with either different or similar labels. In this paper, we call it ungrouped gloss and grouped VSSign annotations. The datasets collected for the most common words of spoken language sincan2020AUTSLLargeScale; kapitanov2023SlovoRussianSign, typical continuous phrases albanie2020BSL1K; albanie2021BBCOxfordBritishSign; adaloglou2021ComprehensiveStudyDeep, or based on an SL dictionary patra2024HierarchicalWindowedGraph primarily have different (ungrouped) labels for similar signs. For instance, according to zuo2023NaturalLanguageAssistedSign, among 2,000 classes of widely used WLASL dataset li2020WLASL, 334 classes form groups of VSSigns. Additional efforts are needed to merge similar VSSigns and assign unique grouped VSSign labels to them.
Among the reviewed datasets, three papers state that VSSigns were grouped. Three papers confirm the presence of ungrouped VSSigns in the presented datasets and propose some techniques to distinguish them (Table 1). To improve VSSigns classification, Hu et al. hu2021GlobalLocalEnhancementNetwork deform a feature map, stretching more informative areas to emphasize non-manual features. Zuo et al. zuo2023NaturalLanguageAssistedSign propose label smoothing depending on their semantic difference and a common latent space for gloss embeddings and vision features to maximize the separability of confusing signs. Other works do not mention any steps to handle VSSigns in the proposed datasets. To our knowledge, no research has examined the impact of VSSigns on the quality of the resulting encoder for downstream tasks. Such a study is one of the topics of this work, using transfer learning to another language as a downstream task example.
@@ 2.4 Multi-Dataset Training
Although researchers complain about insufficient SLR training data gokul2022AddressingResourceScarcity; papadimitriou2023SignLanguageRecognition, the topic of cross-language dataset sharing is poorly exploited. Gokul et al. gokul2022AddressingResourceScarcity implemented a multilingual SLR model for 11 sign languages by simply translating the labels of all the languages into English. The authors themselves admit that their model of combining different languages is primitive and does not make progress for some datasets. Tornay et al. tornay2020TowardsMultilingualSignLanguage train a unified hand movement model using 3 different sign language resources. Then, they optimize the classifier using the target sign language data. However, their cross-lingual model falls short of the monolingual reference. Yin et al. yin2022MLSLTMultilingualSign propose the MLSLT translation network as a single model for multilingual translation. Their work is limited to their rather small datasets and doesn’t address leveraging large SL datasets to improve the model quality. SignCLIP jiang2024SignCLIP utilizes a multilingual sign language dictionary, SpreadTheSign.com, for training. It contains mainly a single sample per concept per language, and the SignCLIP authors translate all concepts into English, similar to gokul2022AddressingResourceScarcity, so knowledge transfer between different sign languages is not investigated. Hu et al. hu2022CollaborativeMultilingualContinuous introduced an additional shared module that learns knowledge from two languages. It improved accuracy for Chinese CSL-Daily zhou2021improving and German Phoenix-14 koller2015ContinuousSignLanguage datasets. Wei et al. wei2023ImprovingContinuousSign also benefit from the joint using the same datasets by creating a gloss translation map based on the visual similarity of signs, rather than their meanings. Authors train the model for the German language using both datasets and replace gloss labels in Chinese videos with German labels using this map, treating Chinese signs as German. As shown below, this mapping method does not give optimal results (see Section 5.3). However, as far as we know, the ability of a model trained on an extensive SL dataset to serve as an encoder for SL tasks for other sign languages has not been explored enough. Such a study, along with the comparison of different approaches to it, is another subject of our work.